NoawieWoawie:

Type the correct answer in each box. Spell all words correctly, and use numerals instead of words for numbers. Use decimal numbers instead of fractions. The equation d = 3t gives the distance, d, in meters that Liam swims with respect to time, t, in seconds. The table gives the different distances in meters that Edgar swam with respect to time in seconds. Time (seconds) Distance (meters) 20 64 40 128 60 192 Assume that Liam and Edgar both swim at a constant rate. On a graph of these relationships, the slope of the line representing the greater unit rate in terms of distance traveled per unit of time would be ___ So, swims at a faster rate than ___

NoawieWoawie:

all you guys have to do is fill in the blanks and explain to me why you got that answer

NoawieWoawie:

are you able to help me, or no?

NoawieWoawie:

sorry, but there should be a blank space after So,

InsatiableSuffering:

Easy. Slope intercept form is your friend here. The formula to find the derivative, or slope, of a table is the following. \[\frac{ f(b)-f(a) }{ b-a }\] Or in other words, plug in corresponding y values of 2 points with their corresponding x values. In this case, it would be \[\frac{ 128-64 }{ 40-20 } \to \frac{ 64 }{ 20 } \to \frac{ 32 }{ 10 } \to \frac{ 16 }{ 5 } =3.2\] Therefore, Edgar swims faster by .2 meters per second.

NoawieWoawie:

oh, okay! thank you!